OpenStudy (anonymous):

Find the geometric mean of 36 and 16.

OpenStudy (asnaseer):

geometric mean of two numbers a and b is defined as:\[\sqrt{a\times b}\]

OpenStudy (anonymous):

\[\sqrt{36X16}\]

OpenStudy (anonymous):

So would i do that?

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

Hint: \(\sqrt{36 \cdot 16}=\sqrt {36} \sqrt{16}\)

OpenStudy (asnaseer):

you should make use of the fact that:\[36=6^2\]and\[16=4^2\]

OpenStudy (anonymous):

so the answer would be 96?

OpenStudy (asnaseer):

how did you get 96?

OpenStudy (anonymous):

the square root of 36X16

OpenStudy (asnaseer):

you can write this as:\[\sqrt{36*16}=\sqrt{36}\times\sqrt{16}\] now, what is \(\sqrt{36}\)? and what is \(\sqrt{16}\)?

OpenStudy (anonymous):

6 and 4

OpenStudy (asnaseer):

correct, so you would get:\[\sqrt{36*16}=\sqrt{36}\times\sqrt{16}=6*4=?\]

OpenStudy (anonymous):

so the geometric mean of 36 and 16 is 24?

OpenStudy (asnaseer):

correct

OpenStudy (anonymous):

all right thx c:

OpenStudy (asnaseer):

yw